# Rankify: A Comprehensive Python Toolkit for Retrieval, Re-Ranking, and Retrieval-Augmented Generation

### v2

@@ Abstract.
Retrieval, re-ranking, and retrieval-augmented generation (RAG) are critical components of modern natural language processing (NLP) applications in information retrieval, question answering, and knowledge-based text generation. However, existing solutions are often fragmented, lacking a unified framework that easily integrates these essential processes. The absence of a standardized implementation, coupled with the complexity of retrieval and re-ranking workflows, makes it challenging for researchers to compare and evaluate different approaches in a consistent environment. While existing toolkits such as Rerankers and RankLLM provide general-purpose reranking pipelines, they often lack the flexibility required for fine-grained experimentation and benchmarking. In response to these challenges, we introduce Rankify, a powerful and modular open-source toolkit designed to unify retrieval, re-ranking, and RAG within a cohesive framework. Rankify supports a wide range of retrieval techniques, including dense and sparse retrievers, while incorporating state-of-the-art re-ranking models to enhance retrieval quality. Additionally, Rankify includes a collection of pre-retrieved datasets to facilitate benchmarking, available at Huggingface [^1] https://huggingface.co/datasets/abdoelsayed/reranking-datasets-light. To encourage adoption and ease of integration, we provide comprehensive documentation [^2] http://rankify.readthedocs.io/, an open-source implementation on GitHub [^3] https://github.com/DataScienceUIBK/rankify, and a PyPI package for effortless installation [^4] https://pypi.org/project/rankify/. By providing a unified and lightweight framework, Rankify allows researchers and practitioners to advance retrieval and re-ranking methodologies while ensuring consistency, scalability, and ease of use.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Information retrievalccs: Information systems Evaluation of retrieval resultsccs: Information systems Presentation of retrieval resultsccs: Information systems Top-k retrieval in databasesccs: Information systems Rank aggregation
@@ 1. Introduction
Information retrieval (IR) systems (Singhal et al., 2001; Chowdhury, 2010) are fundamental to many natural language processing (NLP) applications (Wang et al., 2024; Liu et al., 2024), including question-answering (Khamnuansin et al., 2024; Abdallah et al., 2024b), search engines (Croft et al., 2010; Xiong et al., 2024), and knowledge-based generation (Long et al., 2024). These systems often rely on a two-stage pipeline: a retriever that efficiently identifies a set of candidate documents and a re-ranker that refines these results to maximize relevance to the query (Yates et al., 2021; Abdallah et al., 2024a; Nogueira et al., 2019; Ren et al., 2021). This approach has proven highly effective, with retrieval and re-ranking methods achieving state-of-the-art performance across diverse NLP benchmarks.
Retrievers form the backbone of information retrieval systems by identifying a subset of documents relevant to a user’s query. The retrieval landscape includes sparse, dense, and hybrid methods. Sparse retrievers, such as BM25 (Robertson et al., 1995), rely on exact term matching by representing queries and documents as high-dimensional sparse vectors. These methods are highly effective when query terms closely align with document content but struggle to capture semantic relationships. Dense retrievers, including models like DPR (Yates et al., 2021), BPR (Yamada et al., 2021), ColBERT (Santhanam et al., 2021), BGE (Chen et al., 2023), Contriever (Izacard et al., 2021) and ANCE (Xiong et al., 2021a), overcome this limitation by encoding queries and documents into low-dimensional dense vectors using neural networks, enabling retrieval based on semantic similarity even when lexical overlap is minimal. However, dense retrieval requires significant computational resources for training and inference. Hybrid retrieval methods (Yang et al., 2019; Chen et al., 2022) integrate the strengths of both approaches, combining the precision of sparse retrievers with the semantic understanding of dense models to achieve a more balanced and robust retrieval system.
@@
Toolkit # Datasets # Retrievers # Re-rankers # RAG Methods Modular Design Automatic Evaluation Corpus
FlashRAG (Jin et al., 2024) 32 3 2 12
FastRAG (Abane et al., 2024) 0 1 3 7
AutoRAG (Kim et al., 2024a) 4 2 7 1
LocalRQA (Yu et al., 2024) 0 2 0 0
Rerankers (Clavié, 2024) 0 0 15 0
Rankify 40 7 24 3+
Table 1. Comparison with other Retriever, Re-ranking, and RAG toolkits. Modular Component refers to whether the toolkit is composed of modular components. Automatic Evaluation indicates if the toolkit provides automated evaluation capabilities to assess the performance of various datasets. Corpus Helper shows if the toolkit offers auxiliary tools for processing corpus, including cleaning and chunking.
@@
Re-ranking techniques enhance the initial retrieval results by ensuring the most relevant documents appear at the top. These methods are categorized into three main approaches: (1) Pointwise reranking (Nogueira et al., 2019; Abdallah et al., 2025b, a; Sachan et al., 2022) treats reranking as a regression or classification task, assigning independent relevance scores to each document. (2) Pairwise reranking (Huang and Chen, 2024; Qin et al., 2023; Sinhababu et al., 2024) refines ranking by comparing document pairs and optimizing their relative order based on relevance. However, this approach treats all document pairs equally, sometimes improving lower-ranked results at the expense of top-ranked ones. (3) Listwise reranking (Sun et al., 2023a; Pradeep et al., 2023b; Yoon et al., 2024) considers the entire document list, prompting LLMs with the query and a subset of candidates for reranking. Due to input length limitations, these methods tend to employ a sliding window strategy, progressively refining rankings from back to front.
@@
Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Mozafari et al., 2024; Abdallah and Jatowt, 2023; Gao et al., 2023) enhances language models by integrating retrieval and generation, making them more effective in knowledge-intensive tasks. Rather than relying solely on pre-trained knowledge, RAG dynamically retrieves relevant documents from external sources during inference, incorporating them into the generation process.
@@
The growing complexity and diversity of retrieval, re-ranking, and RAG methods pose significant challenges for benchmarking, reproducibility, and integration. Existing toolkits, such as Pyserini (Lin et al., 2021), Rerankers (Clavié, 2024) and RankLLM (Pradeep et al., 2023a), often lack flexibility, enforce rigid implementations, and require extensive preprocessing, making them less suitable for research-driven experimentation. Additionally, retrieval and re-ranking datasets are scattered across different sources, complicating evaluation and comparison. To address these challenges, we introduce Rankify, an open-source framework that unifies retrieval, re-ranking, and RAG into a modular and extensible ecosystem with its logo shown in Figure 1. Rankify supports diverse retrieval techniques, integrates state-of-the-art re-ranking models, and provides curated pre-retrieved datasets to streamline experimentation. Designed with flexibility, it enables researchers to efficiently build, evaluate, and extend retrieval pipelines while ensuring consistency in benchmarking.
@@
  • Curated retrieval datasets and precomputed embeddings: Rankify provides 40 datasets with 1,000 pre-retrieved documents per query across various domains (QA, dialogue, entity linking, etc.). It also includes pre-computed Wikipedia and MS MARCO corpora for multiple retrievers, eliminating preprocessing overhead.
@@ 2. Related Work
2.1. Retrievers
Retrieval models have evolved significantly, transitioning from traditional sparse retrievers to sophisticated dense and hybrid retrieval techniques. Sparse retrievers, such as BM25 (Robertson et al., 1995), remain widely used due to their strong lexical matching capabilities and generalization. However, they struggle with capturing semantic relationships, leading to the rise of dense retrieval methods (Lee et al., 2019), which leverage pre-trained neural encoders to generate embeddings for queries and documents. Notable advancements, such as DPR (Yates et al., 2021) and ANCE (Xiong et al., 2021a), improve retrieval effectiveness by employing contrastive training with in-batch negatives (Gillick et al., 2019) and hard-negative mining (Karpukhin et al., 2020; Xiong et al., 2021b). Multi-vector approaches like ColBERT (Khattab and Zaharia, 2020) enable fine-grained interaction between queries and documents, albeit at the cost of increased computational overhead. To address the limitations of dense retrievers, hybrid models (Gao et al., 2021; Ma et al., 2022) combine sparse and dense approaches, utilizing both term matching and semantic understanding for enhanced retrieval performance.
Recent research has also explored retrieval enhancements through knowledge distillation, curriculum learning, and pre-training. Distillation techniques transfer knowledge from cross-encoders or teacher retrievers to student models, improving dense retriever effectiveness (Qu et al., 2021; Chen et al., 2021). Curriculum learning strategies, which progressively train retrievers from easy to hard samples (Zeng et al., 2022; Lin et al., 2022a), have been shown to enhance both supervised and zero-shot retrieval effectiveness. Meanwhile, pre-training methods such as contrastive learning (Izacard et al., 2021) and masked autoencoding (Gao and Callan, 2021; Liu and Shao, 2022) have significantly improved the generalization of retrieval models. Additionally, approaches like query augmentation (Mao et al., 2021; Ma et al., 2021) generate synthetic training data to enhance retrieval robustness.
2.2. Re-Ranking Methods
Re-ranking plays a crucial role in refining retrieved document lists to improve ranking quality. Traditional approaches include pointwise learning-to-rank (LTR) methods, where each document is scored independently (Zhu et al., 2021; Cinar and Renders, 2020), and pairwise methods, which optimize the relative ordering of document pairs (Cao et al., 2007; Jia et al., 2021). Listwise approaches, such as LambdaRank and ListNet, directly optimize ranking metrics like Discounted Cumulative Gain (DCG) to enhance top-ranked results (Burges, 2010; Liu et al., 2017; Cao et al., 2007). More recently, deep neural networks and large language models (LLMs) have revolutionized re-ranking, enabling context-aware ranking adjustments. Cross-encoder models, which jointly encode queries and candidate documents, have demonstrated significant improvements in ranking accuracy by capturing fine-grained interactions between them. Additionally, zero-shot re-ranking approaches leverage models like GPT-4 for listwise re-ranking without task-specific training (Achiam et al., 2023; Sun et al., 2023c), while knowledge distillation techniques such as RankZephyr fine-tune smaller models to approximate LLM performance with reduced computational cost (Pradeep et al., 2023b).
Re-ranking is also increasingly integrated into retrieval-augmented generation (RAG) systems, where refining retrieved documents before response generation ensures higher relevance and coherence in generated outputs. These advancements highlight the transition from traditional ranking techniques to sophisticated, model-driven re-ranking approaches that enhance retrieval quality in diverse NLP applications. Recent innovations in re-ranking include leveraging large language models (LLMs) such as GPT-4 (Achiam et al., 2023) for zero-shot and listwise re-ranking (Sun et al., 2023c). Other approaches, such as RankZephyr (Pradeep et al., 2023b), utilize knowledge distillation to fine-tune smaller models based on LLM outputs, enabling efficient and effective re-ranking.
2.3. Retrieval-Augmented Generation (RAG)
Retrieval-Augmented Generation (RAG) combines retrieval and generation to enhance language models in knowledge-intensive tasks. Early RAG systems (Lewis et al., 2020) retrieve context documents relevant to a query and use these as input for generation, enabling models to answer queries beyond their training data. More recent approaches explore the integration of retrieval with both encoder-decoder (Izacard and Grave, 2020) and decoder-only (Raffel et al., 2020) architectures, optimizing performance for a variety of downstream tasks.
One key area of focus in RAG is mitigating hallucinations and improving the factuality of generated responses (Manakul et al., 2023). Techniques such as self-consistency (Wang et al., 2022b) and noisy context filtering (Fang et al., 2024) have been proposed to enhance the reliability of RAG outputs. However, the quality of retrieved documents remains a critical factor in the overall effectiveness of RAG pipelines. Recent works emphasize the importance of combining retrieval with robust re-ranking techniques to ensure that the retrieved context is both relevant and precise (Kamalloo et al., 2023).
2.4. Frameworks
Several frameworks have been developed to support research and development in retrieval and re-ranking systems. rerankers (Clavié, 2024) is a lightweight Python library that provides a unified interface to commonly used re-ranking approaches, simplifying the integration of various methods into retrieval pipelines. Similarly, RankLLM [^5] https://github.com/castorini/rank_llm/tree/main offers a Python toolkit focused on reproducible information retrieval research, emphasizing listwise re-ranking with large language models. Additionally, FlashRAG (Jin et al., 2024) and AutoRAG (Kim et al., 2024b) introduce a modular toolkit designed for efficient retrieval-augmented generation (RAG) research, aiming to enhance flexibility and extensibility in experimental settings. These frameworks, along with others like LangChain (Chase, 2022), LlamaIndex (Liu, 2022), and DSPy (Khattab et al., 2023), provide researchers and practitioners with versatile tools to develop and evaluate retrieval and RAG applications. A comparison of retrieval, re-ranking, and RAG toolkits is presented in Table 1, highlighting the advantages of Rankify in dataset diversity, retriever and re-ranker support, and modularity.
@@ 3. Rankify
To ensure accessibility and reproducibility, Rankify supports both precomputed and real-time retrieval workflows. Users can work with pre-retrieved datasets for benchmarking or perform live document retrieval and re-ranking on large-scale corpora. The framework includes ready-to-use indices for Wikipedia and MS MARCO, eliminating the need for time-consuming indexing. Additionally, it supports different retrievers, giving users the flexibility to test different retrieval paradigms without modifying their pipeline.
Rankify consists of four core modules: Datasets, Retrievers, Re-Rankers, and RAG. The Datasets module provides standardized access to benchmark datasets, including Natural Questions, TriviaQA, and HotpotQA, as well as the ability to create custom datasets. The Retrievers module integrates state-of-the-art retrieval methods, including BM25, DPR, ANCE, BGE, and ColBERT, allowing users to experiment with different retrieval strategies. The Re-Rankers module includes 24 ranking models with 41 sub-methods, supporting pointwise, pairwise, and listwise re-ranking techniques. Finally, the RAG module enables retrieval-augmented generation using models such as LLAMA, GPT, and T5, supporting in-context learning, Fusion-in-Decoder (FiD), and retrieval-augmented latent memory (RALM).
A key advantage of Rankify is its unified API, which abstracts away implementation details, allowing users to experiment with different models and configurations using a simple interface. Figure 2 illustrates the architecture of Rankify, demonstrating how retrieval, ranking, and generation components interact within the pipeline. The framework is designed to be scalable and adaptable, making it suitable for both academic research and real-world applications in search, question answering, and document ranking.
To encourage collaboration and community contributions, Rankify is open-source and actively maintained [^7] https://github.com/abdoelsayed2016/reranking. Extensive documentation [^8] http://rankify.readthedocs.io/ and inline docstrings provide clear guidance on using the framework, ensuring ease of adoption for both beginners and experienced researchers. The next sections explore its capabilities in greater detail, covering retrieval benchmarks, ranking evaluations, and retrieval-augmented generation results.
@@
Figure 3. Summary of datasets used in our Rankify. The figure illustrates the dataset sizes across different categories. The bubble size represents the total number of samples (Train + Dev + Test) for each dataset. Categories include QA, Multi-Hop QA, Long-Form QA, Multiple-Choice, Entity-Linking, Slot Filling, Fact Verification, Dialog Generation, Summarization, and Other specialized datasets. The log scale is used for better visualization of dataset size variations.
Figure 4. Schema of the Dataset class, illustrating the objects used to represent a dataset in Rankify.

### v3

@@ Abstract.
Retrieval, re-ranking, and retrieval-augmented generation (RAG) are critical components of modern applications in information retrieval, question answering, or knowledge-based text generation. However, existing solutions are often fragmented, lacking a unified framework that easily integrates these essential processes. The absence of a standardized implementation, coupled with the complexity of retrieval and re-ranking workflows, makes it challenging for researchers to compare and evaluate different approaches in a consistent environment. While existing toolkits such as Rerankers and RankLLM provide general-purpose reranking pipelines, they often lack the flexibility required for fine-grained experimentation and benchmarking. In response to these challenges, we introduce Rankify, a powerful and modular open-source toolkit designed to unify retrieval, re-ranking, and RAG within a cohesive framework. Rankify supports a wide range of retrieval techniques, including dense and sparse retrievers, while incorporating state-of-the-art re-ranking models to enhance retrieval quality. Additionally, Rankify includes a collection of pre-retrieved datasets to facilitate benchmarking, available at Huggingface [^1] https://huggingface.co/datasets/abdoelsayed/reranking-datasets-light. To encourage adoption and ease of integration, we provide comprehensive documentation [^2] http://rankify.readthedocs.io/, an open-source implementation on GitHub [^3] https://github.com/DataScienceUIBK/rankify, and a PyPI package for easy installation [^4] https://pypi.org/project/rankify/. As a unified and lightweight framework, Rankify allows researchers and practitioners to advance retrieval and re-ranking methodologies while ensuring consistency, scalability, and ease of use.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; July 13–18, 2025; Padova, ITisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Information retrievalccs: Information systems Evaluation of retrieval resultsccs: Information systems Presentation of retrieval resultsccs: Information systems Top-k retrieval in databasesccs: Information systems Rank aggregation
@@ 1. Introduction
Information retrieval (IR) systems (Singhal et al., 2001; Chowdhury, 2010) are fundamental to many applications (Wang et al., 2024; Liu et al., 2024), including question-answering (Khamnuansin et al., 2024; Abdallah et al., 2025b), search engines (Croft et al., 2010; Xiong et al., 2024), and knowledge-based generation (Long et al., 2024). These systems often rely on a two-stage pipeline: a retriever that efficiently identifies a set of candidate documents and a re-ranker that refines these results to maximize relevance to the query (Yates et al., 2021; Nogueira et al., 2019; Ren et al., 2021). This approach has proven highly effective, with retrieval and re-ranking methods achieving state-of-the-art performance across diverse NLP benchmarks.
Table 1. Comparison of Retriever, Re-ranking, and RAG toolkits. Modular Design indicates if the toolkit uses modular components. Automatic Evaluation refers to the availability of built-in evaluation tools. Corpus shows whether the toolkit provides utilities for corpus processing, such as cleaning and chunking. The Combination column represents the total number of possible configurations, calculated as the product of the number of datasets, retrievers, re-rankers, and RAG methods (e.g., for Rankify: 40 × 7 × 24 × 3 = 20,160).
@@
Toolkit # Datasets # Retrievers # Re-rankers # RAG Methods Modular Design Automatic Evaluation Corpus Combination
FlashRAG (Jin et al., 2024) 32 3 2 12 2,304
FastRAG (Abane et al., 2024) 0 1 3 7 21
AutoRAG (Kim et al., 2024) 4 2 7 1 56
LocalRQA (Yu et al., 2024) 0 2 0 0 2
Rerankers (Clavié, 2024) 0 0 15 0 15
CHERCHE (Sourty et al., 2022) 0 7 2 0 14
Rankify 40 7 24 3 20,160
Retrievers form the backbone of information retrieval systems by identifying a subset of documents relevant to a user’s query. The retrieval landscape includes sparse, dense, and hybrid methods. Sparse retrievers, such as BM25 (Robertson et al., 1995), rely on exact term matching by representing queries and documents as high-dimensional sparse vectors. These methods are highly effective when query terms closely align with document content but they struggle to capture semantic relationships. Dense retrievers, including models like DPR (Yates et al., 2021), BPR (Yamada et al., 2021), ColBERT (Santhanam et al., 2021), BGE (Chen et al., 2023), Contriever (Izacard et al., 2021a) and ANCE (Xiong et al., 2021), overcome this limitation by encoding queries and documents into low-dimensional dense vectors using neural networks, enabling retrieval based on semantic similarity even when lexical overlap is minimal. However, dense retrieval requires significant computational resources for training and inference. Hybrid retrieval methods (Yang et al., 2019; Chen et al., 2022) integrate the strengths of both approaches, combining the precision of sparse retrievers with the semantic understanding of dense models to achieve a more balanced and robust retrieval system.
@@
Re-ranking techniques enhance the initial retrieval results by ensuring the most relevant documents appear at the top. These methods are categorized into three main approaches: (1) Pointwise reranking (Nogueira et al., 2019; Abdallah et al., 2025b, a; Sachan et al., 2022) treats reranking as a regression or classification task, assigning independent relevance scores to each document. (2) Pairwise reranking (Huang and Chen, 2024; Qin et al., 2023; Sinhababu et al., 2024) refines ranking by comparing document pairs and optimizing their relative order based on relevance. However, this approach treats all document pairs equally, sometimes improving lower-ranked results at the expense of top-ranked ones. (3) Listwise reranking (Sun et al., 2023; Pradeep et al., 2023b; Yoon et al., 2024) considers the entire document list, prompting LLMs with the query and a subset of candidates for reranking. Due to input length limitations, these methods tend to employ a sliding window strategy, progressively refining rankings from back to front.
@@
Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Gao et al., 2023) enhances language models by integrating retrieval and generation, making them more effective in knowledge-intensive tasks. Rather than relying solely on pre-trained knowledge, RAG dynamically retrieves relevant documents from external sources during inference, incorporating them into the generation process.
@@
The growing complexity and diversity of retrieval, re-ranking, and RAG methods pose significant challenges for benchmarking, reproducibility, and integration. Existing toolkits, such as Pyserini (Lin et al., 2021), Rerankers (Clavié, 2024) and RankLLM (Pradeep et al., 2023a) often lack flexibility, enforce rigid implementations, and require extensive preprocessing, making them less suitable for research-driven experimentation. Additionally, retrieval and re-ranking datasets are scattered across different sources, complicating evaluation and comparison. To address these challenges, we introduce Rankify, an open-source framework that unifies retrieval, re-ranking, and RAG into a modular and extensible ecosystem (logo shown in Figure 1). Rankify supports diverse retrieval techniques, integrates the state-of-the-art re-ranking models, and provides curated pre-retrieved datasets to streamline experimentation. Designed for maximizing flexibility, it enables researchers to efficiently build, evaluate, and extend retrieval pipelines while ensuring consistency in benchmarking.
@@
  • Curated retrieval datasets and precomputed embeddings: Rankify provides 40 datasets, each with 1,000 pre-retrieved documents per query, across various domains (QA, dialogue, entity linking, etc.). It also includes pre-computed Wikipedia and MS MARCO corpora for multiple retrievers, eliminating preprocessing overhead.
@@
Figure 2. An overview of the Rankify pipeline, demonstrating its dual capability for document retrieval. Users can interact with the system either by providing a query to retrieve relevant documents in real-time or by leveraging pre-retrieved datasets already indexed by the framework. The process starts with Pre-Retrieved Datasets & Corpus Indexing, where documents are indexed using both dense (e.g., DPR) and sparse (e.g., BM25) retrievers across corpus like Wikipedia and MS MARCO. Next, in the Retrieval & Re-Ranking stage, the system retrieves candidate documents using dense, or sparse retrieval methods and re-ranks them with pointwise pairwise, or listwise models powered by large language models (LLMs). Finally, the Generator stage applies prompt augmentation and uses models like Fusion-in-Decoder (FiD) to generate accurate and contextually informed answers.
@@ 2. Related Work
Research in information retrieval (IR), re-ranking, and retrieval-augmented generation (RAG) has progressed significantly over the past decade. Traditional retrieval models like BM25 (Robertson et al., 1995) offered robust lexical matching capabilities but struggled with capturing semantic relationships. This limitation led to the development of dense retrieval methods (Lee et al., 2019), which leverage pre-trained neural encoders to represent queries and documents in a shared semantic space. Notable approaches, such as DPR (Yates et al., 2021), ANCE (Xiong et al., 2021), and multi-vector models like ColBERT (Khattab and Zaharia, 2020), have demonstrated substantial improvements in retrieval effectiveness. Hybrid retrievers, combining sparse and dense signals (Gao et al., 2021; Ma et al., 2022), further enhance performance by leveraging both lexical and semantic features. Recent advancements, including knowledge distillation (Qu et al., 2021) and curriculum learning (Zeng et al., 2022), continue to refine retrieval performance across diverse datasets.
Re-ranking methods have also evolved alongside retrieval techniques to improve the ordering of retrieved documents. Traditional pointwise (Zhu et al., 2021) and pairwise (Cao et al., 2007) approaches have given way to listwise methods like LambdaRank and ListNet (Burges, 2010; Liu et al., 2017). Deep neural models, including cross-encoders and transformer-based architectures, have demonstrated remarkable success in re-ranking tasks by capturing complex interactions between queries and documents. Zero-shot and in-context re-ranking with large language models (LLMs), such as GPT-4 (Achiam et al., 2023) and RankT5 (Zhuang et al., 2023), now enable effective ranking adjustments without task-specific training.
Retrieval-Augmented Generation (RAG) has emerged as a powerful paradigm for enhancing generative models in knowledge-intensive tasks (Lewis et al., 2020). By retrieving relevant documents and integrating them into the generative process, RAG systems improve factual accuracy and reduce hallucinations. Techniques like self-consistency (Wang et al., 2022b) and noise filtering (Fang et al., 2024) have been proposed to further improve the reliability of RAG outputs. However, the effectiveness of these systems heavily depends on the quality of retrieved and re-ranked documents, highlighting the need for robust retrieval frameworks.
In response to the growing complexity of IR, re-ranking, and RAG tasks, several frameworks have been introduced. Rerankers (Clavié, 2024) provides a lightweight Python interface for common re-ranking models, while RankLLM focuses on listwise re-ranking with LLMs. Other frameworks like FlashRAG (Jin et al., 2024) and AutoRAG (Kim et al., 2024) offer modular components for RAG experimentation, though they often lack support for diverse datasets and advanced retriever configurations. Tools such as LangChain (Chase, 2022), LlamaIndex (Liu, 2022), and DSPy (Khattab et al., 2023) further contribute to the ecosystem by simplifying model integration and workflow design. A comparison of retrieval, re-ranking, and RAG toolkits is presented in Table 1, highlighting the advantages of Rankify in dataset diversity, retriever and re-ranker support, and modularity.
@@ 3. Rankify
Rankify supports both precomputed and real-time retrieval workflows, allowing users to benchmark with pre-retrieved datasets or perform live document retrieval on large-scale corpora. The framework includes ready-to-use indices for Wikipedia and MS MARCO, eliminating the need for time-consuming indexing.
The architecture of Rankify consists of four core modules:
  • Datasets: Provides standardized access to datasets like Natural Questions, TriviaQA, and HotpotQA, with support for custom dataset creation.
  • Retrievers: Integrates diverse retrieval methods such as BM25, DPR, ANCE, BGE, and ColBERT, enabling flexible retrieval strategies.
  • Re-Rankers: Includes 24 models with 41 sub-methods, supporting pointwise, pairwise, and listwise re-ranking.
  • RAG: Facilitates retrieval-augmented generation with models like LLAMA, GPT, and T5, supporting in-context learning, Fusion-in-Decoder (FiD), and Incontext RALM.
A key advantage of Rankify is its unified API, which abstracts implementation details and simplifies experimentation across different models. Figure 2 and  3 illustrates the interactions between these modules within the pipeline. The framework is scalable, adaptable, and suited for both research and real-world applications like search engines and question-answering systems. Rankify is open-source, actively maintained [^6] https://github.com/abdoelsayed2016/reranking, and accompanied by comprehensive documentation [^7] http://rankify.readthedocs.io/ to support easy adoption for both beginners and experts.
@@
Table 2. Summary of datasets used in Rankify. Categories include QA, Multi-Hop QA, Long-Form QA, Multiple-Choice, Entity-Linking, Slot Filling, Fact Verification, Dialog Generation, Summarization, and Other specialized datasets.
Task Dataset Name # Train # Val # Test
QA NQ (Kwiatkowski et al., 2019) 79,168 8,757 3,610
TriviaQA (Joshi et al., 2017) 78,785 8,837 11,313
WebQ (Berant et al., 2013) 3,778 - 2,032
SQuAD (Rajpurkar et al., 2016) 87,599 10,570 -
NarrativeQA (s Koˇciský et al., 2018) 32,747 3,461 10,557
MSMARCO-QA (Nguyen et al., 2017) 808,731 101,093 -
PopQA (Mallen et al., 2022) - - 14,267
SIQA (Sap et al., 2019) 33,410 1,954 -
Fermi (Kalyan et al., 2021) 8,000 1,000 1,000
WikiQA (Yang et al., 2015) 20,360 2,733 6,165
AmbigQA (Min et al., 2020; Kwiatkowski et al., 2019) 10,036 2,002 -
CommenseQA (Talmor et al., 2019) 9,741 1,221 -
PIQA (Bisk et al., 2019) 16,113 1,838 -
BoolQ (Clark et al., 2019) 9,427 3,270 -
Multi-Hop QA 2WikiMultiHopQA (Ho et al., 2020) 15,000 12,576 -
Bamboogle (Press et al., 2023) - - 125
Musique (Trivedi et al., 2022) 19,938 2,417 -
HotpotQA (Yang et al., 2018) 90,447 7,405 -
Temp QA ArchivalQA (Wang et al., 2022a) 384,426 48,304 48760
ChroniclingQA (Piryani et al., 2024) 385,629 21,739 21,735
Long-Form QA ELI5 (Fan et al., 2019) 272,634 1,507 -
ASQA (Stelmakh et al., 2022) 4,353 948 -
Multiple-Choice MMLU (Hendrycks et al., 2021b, a) 99,842 1,531 14,042
TruthfulQA (Lin et al., 2022) - 817 -
HellaSwag (Zellers et al., 2019) 39,905 10,042 -
ARC (Clark et al., 2018) 3,370 869 3,548
OpenBookQA (Mihaylov et al., 2018) 4,957 500 500
Entity-linking WNED (Tedeschi et al., 2021; Petroni et al., 2021) - 8,995 -
AIDA CoNLL-YAGO (Hoffart et al., 2011; Petroni et al., 2021) 18,395 4,784 -
Slot filling Zero-shot RE (Levy et al., 2017; Petroni et al., 2021) 147,909 3,724 -
T-REx (ElSahar et al., 2018; Petroni et al., 2021) 2,284,168 5,000 -
Dialog Generation WOW (Dinan et al., 2019; Petroni et al., 2021) 63,734 3,054 -
Fact Verification FEVER (Thorne et al., 2018; Petroni et al., 2021) 104,966 10,444 -
Open-domain Summarization WikiAsp (Hayashi et al., 2020) 300,636 37,046 37,368